# The Digital Ecosystem of Beliefs: does evolution favour AI over humans?

### v2

@@ Appendix F The effect of truthfulness on beliefs
The content that one creates may not always reflect one’s true beliefs. Sometimes agents may send messages that disagree with their beliefs to make other agents come to their side (i.e. make the belief weight for that agent larger). Assuming all agents behave equally truthfully or deceptive, Hypothesis 5 states that truthfulness (high $\lambda$) will lead to agents being less subject to manipulation (i.e. lower Human Belief 0). To evaluate the hypothesis, we consider the $\lambda$ parameter, which affects the belief penalty in the fitness function (see Eq. 8), as the truthfulness and manipulate it in $\{0.01,0.1,1.0,10.0\}$. To assess propaganda effectiveness, only the All Zero conditions are included.

### v3

@@ Appendix F The effect of truthfulness on beliefs
The content that one creates may not always reflect one’s true beliefs. Sometimes agents may send messages that disagree with their beliefs to make other agents come to their side (i.e. make the belief weight for that agent larger). Assuming all agents behave equally truthfully or deceptive, Hypothesis 5 states that truthfulness (high $\lambda$) will lead to agents being less subject to manipulation (i.e. lower Human Belief 0). To evaluate the hypothesis, we consider the $\lambda$ parameter, which affects the belief penalty in the fitness function (see Eq. 8 in the main text), as the truthfulness and manipulate it in $\{0.01,0.1,1.0,10.0\}$. To assess propaganda effectiveness, only the All Zero conditions are included.